In a recent game the mean score was 200 and the standard deviation was 25.
(a) Find the z-score of Jacob who scored 120.
(b) Find the z-score of Layla who scored 250.
(c) If Nancy had a z-score of 1.5, what was her score?

Answer by stanbon(75887)

n a recent game the mean score was 200 and the standard deviation was 25.
(a) Find the z-score of Jacob who scored 120.::
Z(120) = (120-200)/25 = -80/25 = -3.2
------------------------------------------------
(b) Find the z-score of Layla who scored 250.
Z(250) = (250-200)/25 = 2
-----------------------------
(c) If Nancy had a z-score of 1.5, what was her score?
X = Z*S + Mean
X = 1.5*25 + 200
X = 237.50
------------
Cheers,
Stan H.
-------